# A.S.E: A Repository-Level Benchmark for Evaluating Security in AI-Generated Code

### v2

@@ 5 Conclusion
This paper addressed the critical need for realistic evaluation of LLM-generated code security, as existing benchmarks often use isolated snippets and lack reproducible, context-aware methods. We introduced A.S.E (AI Code Generation Security Evaluation), a benchmark built from real-world repositories with documented CVEs. By preserving full project context and using a reproducible, containerized framework with expert-defined rules, A.S.E reliably assesses code security, build quality, and stability, directly linking them to the model’s input. Our evaluation of leading LLMs yielded three key findings: (1) While Claude-3.7-Sonnet leads in overall performance, (2) the security gap to open-source models is narrow, with Qwen3-235B-A22B-Instruct achieving the top security score. (3) Crucially, concise "fast-thinking" decoding strategies consistently outperform complex "slow-thinking" reasoning for security repairs. These results have significant implications, suggesting prompting strategy is as critical as model choice and that open-source models are highly competitive in security. A.S.E provides a robust foundation for future research into developing more secure and reliable LLMs for software engineering.
6 Data Availability
To facilitate further research and practical adoption, we have made both the source code of the A.S.E evaluation framework and the full benchmark dataset publicly available, which can be accessed at https://github.com/Tencent/AICGSecEval. In addition, we provide comprehensive evaluation results for all tested models, which are continuously updated and maintained on the project website https://aicgseceval.tencent.com/home. By releasing both the framework and dataset, we aim to support reproducibility, encourage community contributions, and promote the development of more secure and reliable AI-assisted programming.

### v3

@@ 6 Conclusion
In this work, we introduced A.S.E (AI Code Generation Security Evaluation), the first repository-level benchmark dedicated to evaluating the security of AI-generated code. Unlike existing security-oriented benchmarks, A.S.E is constructed from real-world projects with documented CVEs, incorporates repository-level context extraction, and integrates customized static vulnerability detection rules to provide accurate and reproducible evaluations. Through extensive experiments on $26$ commercial and open-source LLMs, we demonstrated that while current models exhibit strong performance in functional correctness, they continue to face substantial challenges in secure code generation. Repository-level tasks further expose their limitations. These observations provide valuable insights into the current state of AI code generation. They not only help developers choose appropriate models and prompting strategies for practical tasks but also offer a foundation for refining LLMs toward generating secure, efficient, and reliable code in real-world environments. Beyond evaluation, A.S.E contributes to the ecosystem by supporting model comparison, prompt engineering, enterprise adoption pipelines, and educational use cases. While it does not yet capture the full diversity of software engineering practices, A.S.E marks a substantial advance beyond prior benchmarks and represents an important step toward security-aware evaluation of AI-assisted programming.
Data Availability
To facilitate further research and practical adoption, we have made both the source code of the A.S.E evaluation framework and the full benchmark dataset publicly available, which can be accessed at https://github.com/Tencent/AICGSecEval. In addition, we will provide comprehensive evaluation results for all tested models, which are continuously updated and maintained on the project website. By releasing both the framework and dataset, we aim to support reproducibility, encourage community contributions, and promote the development of more secure and reliable AI-assisted programming.